# Consistency of variational inference for Besov priors in non-linear inverse problems

### v1

@@ Abstract.
This study investigates the variational posterior convergence rates of inverse problems for partial differential equations (PDEs) with parameters in Besov spaces $B_{pp}^{\alpha}$ ($p\geq 1$) which are modeled naturally in a Bayesian manner using Besov priors constructed via random wavelet expansions with $p$-exponentially distributed coefficients. Departing from exact Bayesian inference, variational inference transforms the inference problem into an optimization problem by introducing variational sets. Building on a refined “prior mass and testing” framework, we derive general conditions on PDE operators and guarantee that variational posteriors achieve convergence rates matching those of the exact posterior under widely adopted variational families (Besov-type measures or mean-field families). Moreover, our results achieve minimax-optimal rates over $B^{\alpha}_{pp}$ classes, significantly outperforming the suboptimal rates of Gaussian priors (by a polynomial factor). As specific examples, two typical nonlinear inverse problems, the Darcy flow problems and the inverse potential problem for a subdiffusion equation, are investigated to validate our theory. Besides, we show that our convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal.
@@ 1. Introduction
To extract information from the posterior probability distribution, sampling methods such as Markov chain Monte Carlo (MCMC) methods are frequently utilized. While MCMC is highly efficient and theoretically sound as a sampling method, its computational cost can become excessive for PDE-constrained likelihood evaluations [13]. This bottleneck motivates variational Bayesian inference as a popular alternative. By minimizing the Kullback-Leibler (KL) divergence between a tractable family $\mathcal{Q}$ and the true posterior $\Pi(\cdot|D_{N})$, variational inference seeks to identify the variational posterior $\hat{Q}$, the closest approximation to $\Pi(\cdot|D_{N})$. The structure of variational sets $\mathcal{Q}$ often allows variational inference methods to achieve comparable accuracy to MCMC with orders-of-magnitude speed improvements [47]. Their optimization-based architecture particularly excels in large-scale inverse problems involving computationally intensive likelihood functions (see e.g. [8, 31]). This approach is increasingly favored in Bayesian inverse problems, as illustrated by recent studies [40, 25] and their references.
@@
$\varepsilon_{N}^{2}+\frac{1}{N}\mathop{\inf}_{Q\in\mathcal{Q}}P_{0}^{(N)}D(Q\|\Pi_{N}(\cdot|D_{N})).$
The first term, $\varepsilon_{N}^{2}$, represents the convergence rate of the posterior $\Pi_{N}(\cdot|D_{N})$ and is ordered as $N^{-a}$ for some $a>0$. The second term represents the variational approximation error arising from the data-generating process $P_{0}^{(N)}$, which is induced by the true parameter. When the variational set $\mathcal{Q}$ comprises Besov-type measures or a mean-field variational class, we demonstrate that the variational approximation error is dominated by $\varepsilon_{N}^{2}$ (up to a logarithmic factor). Therefore, this implies that the convergence rates of the variational posterior distributions for nonlinear inverse problems can attain the same convergence rates as the posterior distributions (up to a logarithmic factor). Besides, in a setting with direct observations ($\mathcal{G}\equivId$), our convergence rate achieves the minimax rate (up to a logarithmic factor) over Besov function classes $B_{pp}^{\alpha}$ for $p\in[1,2)$, while Gaussian process priors can only achieve polynomially slower convergence rates (see Section 4 of [5] for details). In Section 4, we derive the convergence rates of the variational posterior for the Darcy flow problem and the inverse potential problem for a subdiffusion equation. We also show that the convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal.
@@
Variational Bayes inference has seen extensive use across various fields. Recently, theoretical results start to surface. In [46], Wang and Blei established Bernstein–von Mises type of results for parametric models. For other related references on theories on parametric variational Bayes inference, refer to [8]. Regarding nonparametric and high-dimensional models that are more relevant to non-linear inverse problems, recent researches [47, 7] investigated variational approximation to tempered posteriors. In the context of variational Bayesian inference for usual posterior distributions, studies [49, 39] achieved results comparable to those obtained for tempered posteriors. When it comes to inverse problems, some algorithmic developments can be found in [25, 26, 17]. For theoretical results of linear inverse problems, Randrianarisoa and Szabo [41] employed the inducing variable method to derive the contraction rates of variational posterior with Gaussian process priors around the true function. Concerning non-linear inverse problems, study [50] extends these theoretical results of variational posterior to non-linear inverse problem settings with Gaussian priors.
@@
The structure of this paper is as follows. In Section 2, we begin by introducing the notations and statistical settings for the inverse regression model. We also define Besov-type priors on $L^{2}$. In Section 3, we establish a theory that guarantees convergence rates of variational Bayesian inference with Besov-type priors for nonlinear inverse problems. In Section 4, we demonstrate the application of our theory to the Darcy flow problem and the inverse potential problem for a subdiffusion equation, showing that the convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal (up to a logarithmic factor). Proofs for results of inverse problems and properties of Besov priors are given in Sections 5 and 6 respectively.
@@
We repeatedly use the inequalities from [44, Remark 1 on p.143 and Theorem 2.8.3]
@@ 3.1. Conditions for the forward map
The linear subspace $(\mathcal{R},\|\cdot\|_{\mathcal{R}})$ in Conditions 3.1 and 3.2 is usually chosen as the support of Besov priors, such as the support of $\Pi$ defined in (2.8), $(B_{pp}^{b}(\mathcal{Z}),{\|\cdot\|_{B_{pp}^{b}(\mathcal{Z})}})$ for $b<\alpha-\frac{d}{p}$. These regularity and stability conditions on the forward map $\mathcal{G}$ require explicit estimates on growth rates of coefficients. We will verify these conditions for the inverse problems related to the Darcy flow equation and a subdiffusion equation in Section 4.
@@ 3.2. Results for rescaled Besov priors
We estimate $\mathop{\inf}_{Q\in\mathcal{Q}}R(Q)$ in the next theorem. Define a p-exponential mean-field family $\tilde{\mathcal{Q}}_{E}^{J}$ as
$\Bigg{\{}Q=\mathop{\bigotimes}_{l=-1}^{+\infty}\mathop{\bigotimes}_{r\in R_{l}}Exp(p;a_{lr},b_{lr}):a_{lr}\in\mathbb{R},b_{lr}\geq 0\Bigg{\}}.$ (3.7)
@@
The extra logarithmic factor arises in convergence rates due to the product form of the measures in $\mathcal{Q}_{E}$. We note that Theorem 3.5 maintains its convergence rates when we extend our variational set from $\mathcal{Q}_{G}$ to $\mathcal{Q}_{MF}$, which is defined by the push-forward of a standard mean-field family. The set $\mathcal{Q}_{MF}$ is explicitly given by
$\mathcal{Q}_{MF}=\Bigg{\{}Q=\tilde{Q}\circ\Psi^{-1}:d\tilde{Q}(\tilde{\theta})=\prod_{l=-1}^{+\infty}\prod_{r\in R_{l}}d\tilde{Q}_{lr}(\tilde{\theta}_{lr})\Bigg{\}},$ (3.13)
@@ 3.3. Results for high-dimensional Besov priors
Although Theorem 3.9 mainly addresses the case $p\geq 2$, convergence rates under high-dimensional priors remain attainable for $p\in[1,2)$ (see the proof of Theorem 3.9). This extension requires strengthening the regularity requirement for $\theta_{0}$, specifically imposing a more critical regularity constraint $\alpha_{0}\geq\alpha+d/p-d/2$. Failure to meet this condition creates a discrepancy between the $L^{2}$-approximation error of $P_{J}(\theta_{0})$ relative to $\theta_{0}$ and the convergence rates of the variational posterior, arising from the absence of square-integrability (see Lemma 5.2).
@@ 4. Contraction rate for two typical inverse problems
In this section, we apply Theorem 3.5 to the Darcy flow problem and the inverse potential problem for a subdiffusion equation. In order to verify Conditions 3.1 and 3.2 for these problems, we use link functions satisfying specific properties as in [38].
@@ 4.1. Darcy flow problem
We note that for the “PDE-constrained regression” problem of recovering $u_{f_{0}}$ in “prediction” loss, the convergence rate obtained in (4.4) can be shown to be minimax optimal (up to a logarithmic factor) [16, Section 2.3.2]. For a smooth truth $f_{0}$, both of the rates obtained in (4.4) and (4.5) approach the optimal rate $N^{-1/2}$ of finite-dimensional models as $\alpha\rightarrow+\infty$. In particular, our variational reconstruction rate of $f_{0}$ matches the exact posterior contraction rates established in [36] for Gaussian process priors. However, the optimal reconstruction rate for the Darcy flow problem with general Besov regularity of the truth $f_{0}$ remains to be studied for future research.
@@ 4.2. Inverse potential problem for a subdiffusion equation
It is observed that for the “PDE-constrained regression” problem of recovering $u_{q_{0}}(T)$ in “prediction” loss, the convergence rate found in (4.10) will be demonstrated to be minimax optimal (up to a logarithmic factor), as evidenced by Theorem 4.5 given below. The reconstruction rate of $f_{0}$ obtained in (4.11) increases with $p$ because $s/(2+s)$ can have a higher value. However, the optimal reconstruction rate with general Besov regularity of the truth $f_{0}$ remains to be studied for future research. For a smooth truth $q_{0}$, the rates obtained in (4.10) and (4.11) both approach the optimal rate $N^{-1/2}$ of finite-dimensional models as $\alpha\rightarrow+\infty$.
@@
The convergence rates with high-dimensional priors can also be obtained for this inverse problem using Theorem 3.9, which improves the rates of recovering $f_{0}$, since the support of truncated priors has higher regularity. The convergence rates of recovering $f_{0}$ with high-dimensional priors can achieve the minimax optimal rate when $p=2$ (see Theorem 4.4 in [50]).
@@
For $M_{0}>1$, $p\in[1,2]$, $\alpha\in\mathbb{N}$, $q\in\mathcal{F}^{\alpha}_{p,M_{0}}$, consider the solution $u_{q}(t)$ of the problem (4.6). Then there exist fixed $T_{0}>0$ and a finite constant $C>0$ such that for $N$ large enough, the terminal time $T\geq T_{0}$ and any $\eta$,

### v2

@@ Abstract.
This study investigates the variational posterior convergence rates of inverse problems for partial differential equations (PDEs) with parameters in Besov spaces $B_{pp}^{\alpha}$ ($p\geq 1$) which are modeled naturally in a Bayesian manner using Besov priors constructed via random wavelet expansions with $p$-exponentially distributed coefficients. Departing from exact Bayesian inference, variational inference transforms the inference problem into an optimization problem by introducing variational sets. Building on a refined “prior mass and testing” framework, we derive general conditions on PDE operators and guarantee that variational posteriors achieve convergence rates matching those of the true posterior under widely adopted variational families (Besov-type measures or mean-field families). Moreover, our results achieve minimax-optimal rates over $B^{\alpha}_{pp}$ classes, significantly outperforming the suboptimal rates of Gaussian priors (by a polynomial factor). As specific examples, two typical nonlinear inverse problems, the Darcy flow problems and the inverse potential problem for a subdiffusion equation, are investigated to validate our theory. Besides, we show that our convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal.
@@ 1. Introduction
To extract information from the posterior probability distribution, sampling methods such as Markov chain Monte Carlo (MCMC) methods are frequently utilized [49]. While MCMC is highly efficient and theoretically sound as a sampling method, its computational cost can become excessive for PDE-constrained likelihood evaluations [16]. This bottleneck motivates variational Bayesian inference as a popular alternative. By minimizing the Kullback-Leibler (KL) divergence between a tractable family $\mathcal{Q}$ and the true posterior $\Pi(\cdot|D_{N})$, variational inference seeks to identify the variational posterior $\hat{Q}$, the closest approximation to $\Pi(\cdot|D_{N})$. The structure of variational sets $\mathcal{Q}$ often allows variational inference methods to achieve comparable accuracy to MCMC with orders-of-magnitude speed improvements [54]. Their optimization-based architecture particularly excels in large-scale inverse problems involving computationally intensive likelihood functions (see e.g. [9, 35]). This approach is increasingly favored in Bayesian inverse problems, as illustrated by recent studies [45, 28] and their references.
@@
$\varepsilon_{N}^{2}+\frac{1}{N}\mathop{\inf}_{Q\in\mathcal{Q}}P_{0}^{(N)}D(Q\|\Pi_{N}(\cdot|D_{N})),$
which originates from the general contraction theory of variational posterior [56]. The first term, $\varepsilon_{N}^{2}$, represents the convergence rate of the posterior $\Pi_{N}(\cdot|D_{N})$ and is ordered as $N^{-a}$ for some $a>0$. The second term represents the variational approximation error arising from the data-generating process $P_{0}^{(N)}$, which is induced by the true parameter. When the variational set $\mathcal{Q}$ comprises Besov-type measures or a mean-field variational class, we demonstrate that the variational approximation error is dominated by $\varepsilon_{N}^{2}$ (up to a logarithmic factor). Therefore, this implies that the convergence rates of the variational posterior distributions for nonlinear inverse problems can attain the same convergence rates as the posterior distributions (up to a logarithmic factor). Besides, in a setting with direct observations ($\mathcal{G}\equivId$), our convergence rate achieves the minimax rate (up to a logarithmic factor) over Besov function classes $B_{pp}^{\alpha}$ for $p\in[1,2)$, while Gaussian process priors can only achieve polynomially slower convergence rates (see Section 4 of [5] for details). In Section 4, we derive the convergence rates of the variational posterior for the Darcy flow problem and the inverse potential problem for a subdiffusion equation. We also show that the convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal.
@@
Variational Bayes inference has seen extensive use across various fields. Recently, theoretical results start to surface. In [53], Wang and Blei established Bernstein–von Mises type of results for parametric models. For other related references on theories on parametric variational Bayes inference, refer to [9]. Regarding nonparametric and high-dimensional models that are more relevant to non-linear inverse problems, recent researches [54, 7] investigated variational approximation to tempered posteriors. In the context of variational Bayesian inference for usual posterior distributions, studies [56, 44] achieved results comparable to those obtained for tempered posteriors. When it comes to inverse problems, some algorithmic developments can be found in [28, 29, 20]. For theoretical results of linear inverse problems, Randrianarisoa and Szabó [46] employed the inducing variable method to derive the contraction rates of variational posterior with Gaussian process priors around the true function. Concerning non-linear inverse problems, study [57] extends these theoretical results of variational posterior to non-linear inverse problem settings with Gaussian priors.
Our present work extends the prior framework in [57] from Gaussian priors to the broader class of Besov priors generated by $p$-exponential wavelet coefficients. From our previous work, we adopt the same statistical model with Gaussian noise, the general conditions on the forward map $\mathcal{G}$ (local Lipschitz continuity and conditional stability, now formulated in Besov norms), and the overall proof strategy, which combines variational contraction theory from [56] with posterior concentration techniques for nonlinear inverse problems from [42].
The new contributions relative to [57] are as follows. We establish new concentration results for Besov priors. For variational inference, we introduce a novel $p$-exponential mean‑field family $\mathcal{Q}_{E}$ tailored to the Besov prior, and prove that its approximation error is controlled by the contraction rate of the true posterior. We also apply the theory to the Darcy flow and subdiffusion inverse problems, obtaining minimax optimal contraction rates in Besov spaces $B_{pp}^{\alpha}$ for the “PDE‑constrained regression problem”, whereas the rates are polynomially slower under Gaussian priors.
@@
The structure of this paper is as follows. In Section 2, we begin by introducing the notations and statistical settings for the inverse regression model. We also define Besov-type priors on $L^{2}$. In Section 3, we establish a theory that guarantees convergence rates of variational Bayesian inference with Besov-type priors for nonlinear inverse problems. In Section 4, we demonstrate the application of our theory to the Darcy flow problem and the inverse potential problem for a subdiffusion equation, showing that the convergence rates of “prediction” loss for these “PDE-constrained regression problems” are minimax optimal (up to a logarithmic factor). Proofs for results of inverse problems and properties of Besov priors are given in Sections 6 and 7 respectively.
@@
We repeatedly use the inequalities from [51, Remark 1 on p.143 and Theorem 2.8.3]
@@ 3.1. Conditions for the forward map
The linear subspace $(\mathcal{R},\|\cdot\|_{\mathcal{R}})$ in Conditions 3.1 and 3.2 is usually chosen as the support of Besov priors, such as $(B_{pp}^{b}(\mathcal{Z}),{\|\cdot\|_{B_{pp}^{b}(\mathcal{Z})}})$ for $b<\alpha-\frac{d}{p}$. These regularity and stability conditions on the forward map $\mathcal{G}$ require explicit estimates on growth rates of coefficients. We will verify these conditions for the inverse problems related to the Darcy flow equation and a subdiffusion equation in Section 4.
@@ 3.2. Results for rescaled Besov priors
We estimate $\mathop{\inf}_{Q\in\mathcal{Q}}R(Q)$ in the next theorem. Define a p-exponential mean-field family $\tilde{\mathcal{Q}}_{E}$ as
$\Bigg\{{\color[rgb]{0,0,0}\tilde{Q}}=\mathop{\bigotimes}_{l=-1}^{+\infty}\mathop{\bigotimes}_{r\in R_{l}}Exp(p;a_{lr},b_{lr}):a_{lr}\in\mathbb{R},b_{lr}\geq 0\Bigg\}.$ (3.7)
@@
The extra logarithmic factor arises in convergence rates due to the product form of the measures in $\mathcal{Q}_{E}$. For richer variational families, this logarithmic factor could be eliminated. For example, we define the variational family $\mathcal{Q}_{E}^{b}$ to be
[…]
where $Q=\bar{Q}|_{A}$ means that for any measurable set $B$, $Q(B)=\bar{Q}(B\cap A)/\bar{Q}(A)$, that is, the restriction of the measure $\bar{Q}$ on a bounded set $A$. It has been proved that with this variational family, we can eliminate the $\log N$ factor when we estimate the upper bound of $\gamma_{N}^{2}$ (see Theorem 3.5 in [57] and its proof). Besides, we note that (3.11) and (3.12) maintain their convergence rates when we extend our variational set from $\mathcal{Q}_{E}$ to $\mathcal{Q}_{MF}$, which is defined by the push-forward of a standard mean-field family. The set $\mathcal{Q}_{MF}$ is explicitly given by
$\mathcal{Q}_{MF}=\Bigg\{Q=\tilde{Q}\circ\Psi^{-1}:d\tilde{Q}(\tilde{\theta})=\prod_{l=-1}^{+\infty}\prod_{r\in R_{l}}d\tilde{Q}_{lr}(\tilde{\theta}_{lr})\Bigg\},$
@@ 3.3. Results for high-dimensional Besov priors
Although Theorem 3.9 mainly addresses the case $p\geq 2$, convergence rates under high-dimensional priors remain attainable for $p\in[1,2)$ (see the proof of Theorem 3.9). This extension requires strengthening the regularity requirement for $\theta_{0}$, specifically imposing a more critical regularity constraint $\alpha_{0}\geq\alpha+d/p-d/2$. Failure to meet this condition creates a discrepancy between the $L^{2}$-approximation error of $P_{J}(\theta_{0})$ for $\theta_{0}$ and the convergence rates of the variational posterior, arising from the absence of square-integrability (see Lemma 6.2).
@@ 4. Contraction rate for two typical inverse problems
In this section, we apply Theorem 3.5 to the Darcy flow problem and the inverse potential problem for a subdiffusion equation. In order to verify Conditions 3.1 and 3.2 for these problems, we use link functions satisfying specific properties as in [40]. We note that the following results on specific nonlinear inverse problems can be directly extended to the case where $p>2$ using Theorem 3.9 with the truncated Besov prior.
@@ 4.1. Darcy flow problem
We note that for the “PDE-constrained regression” problem of recovering $u_{f_{0}}$ in “prediction” loss, the convergence rate obtained in (4.4) can be shown to be minimax optimal (up to a logarithmic factor) [19, Section 2.3.2]. For a smooth truth $f_{0}$, both of the rates obtained in (4.4) and (4.5) approach the optimal rate $N^{-1/2}$ of finite-dimensional models as $\alpha\rightarrow+\infty$. In particular, our variational reconstruction rate of $f_{0}$ matches the true posterior contraction rates established in [42] for Gaussian process priors. However, the optimal reconstruction rate for the Darcy flow problem with general Besov regularity of the truth $f_{0}$ remains to be studied for future research.
@@ 4.2. Inverse potential problem for a subdiffusion equation
It is observed that for the “PDE-constrained regression” problem of recovering $u_{q_{0}}(T)$ in “prediction” loss, the convergence rate found in (4.10) will be demonstrated to be minimax optimal (up to a logarithmic factor), as evidenced by Theorem 4.5 given below. The reconstruction rate of $q_{0}$ obtained in (4.11) increases with $p$ because $s/(2+s)$ can have a higher value. However, the optimal reconstruction rate with general Besov regularity of the truth $q_{0}$ remains to be studied for future research. For a smooth truth $q_{0}$, the rates obtained in (4.10) and (4.11) both approach the optimal rate $N^{-1/2}$ of finite-dimensional models as $\alpha\rightarrow+\infty$.
@@
The convergence rates with high-dimensional priors can also be obtained for this inverse problem using Theorem 3.9, which improves the rates of recovering $q_{0}$, since the support of truncated priors has higher regularity. The convergence rates of recovering $q_{0}$ with high-dimensional priors can achieve the minimax optimal rate when $p=2$ (see Theorem 4.4 in [57]).
@@
For $M_{0}>1$, $p\in[1,2]$, $\alpha\in\mathbb{N}$, $q\in\mathcal{F}^{\alpha}_{p,M_{0}}$ , consider the solution $u_{q}(t)$ of the problem (4.6). Then there exist fixed $T_{0}>0$ and a finite constant $C>0$ such that for $N$ large enough, the terminal time $T\geq T_{0}$ and any $\eta>0$,
@@
5. Conclusion
5.1. Summary of findings
We establish convergence rates for variational posteriors in nonlinear inverse problems under Besov priors induced by $p$-exponential wavelet coefficients, thereby extending the Gaussian-prior framework to this broader class. Under suitable regularity and stability conditions on the forward map, the variational posterior contracts toward the true parameter at the same rate as the posterior. We also develop new technical tools for Besov priors to control prior mass near the truth. Furthermore, we propose a $p$-exponential mean-field variational family and construct a distribution achieving approximation error $\gamma_{N}^{2}\lesssim\varepsilon_{N}^{2}\log N$, such that the variational posterior inherits the posterior contraction rate. The theory is applied to the Darcy flow and subdiffusion inverse problems, yielding minimax-optimal rates for the “PDE-constrained regression problems”.
5.2. Limitations and future directions
Despite these advances, several limitations of the current work and directions for future research should be acknowledged. By addressing the limitations outlined below and pursuing the proposed directions, we hope to further advance the theory of variational inference for Bayesian inverse problems.
1. Assumptions on the forward map. Conditions 3.1 and 3.2 require polynomial growth bounds on the uniform norm, Lipschitz constant, and stability constant of $\mathcal{G}$. For the PDEs considered, these bounds are verified using specially designed link functions; the common exponential link function $e^{x}$ would instead yield exponential growth. Severely ill‑posed problems (often with logarithmic stability) may also lead to exponential bounds. One could enforce constant bounds by restricting the prior to a ball in Besov space [15], but this requires prior knowledge of the Besov norm of the true parameter, which may therefore be impractical. Relaxing these conditions remains an open problem.
2. Regularity requirements on the truth. The contraction results require the true parameter $\theta_{0}$ to belong to $B_{pp}^{\alpha}(\mathcal{Z})$, with $\alpha$ sufficiently large relative to $d$ and $p$, yet parameters with low regularity are common in practice. A recent study reduces the regularity requirement at the cost of slower contraction rates, so such a result may not be minimax optimal [15]. How to bridge this gap remains an open question.
3. Adaptive variational inference. The rescaled Besov prior uses a scaling $\rho=(N\varepsilon_{N}^{2})^{-1/p}$ with $\varepsilon_{N}=N^{-\frac{\alpha+\kappa}{2\alpha+2\kappa+d}}$, so the prior depends explicitly on sample size and on the regularity $\alpha$ and smoothing index $\kappa$. We hope to adaptively choose $\alpha$ and preserve the optimal contraction rate . Exact Bayesian methods have attained optimal adaptive contraction rates for linear problems [50, 33], whereas this remains an open problem for general nonlinear inverse problems. Variational Bayes is gradually developing an adaptive theory for linear problems, as demonstrated by recent advances in hierarchical variational Bayes [43]. However, this remains an open problem for general nonlinear inverse problems.
4. General Banach spaces. A recent study has developed the contraction theory of the Bayesian posterior for linear inverse problems on general Banach spaces [11]. It may also be possible to extend our theory to certain abstract evolution equations defined on general Banach spaces.